# Double Robustness of Local Projections

### v1

@@ 4.4 Bias-aware inference
We conclude that, while bias-aware VAR inference is possible in theory, in practice the gains relative to the simpler LP interval are small at best, unless we put an extremely tight bound on the noise-to-signal ratio.
@@ 5 Simulations
In this section we show that our asymptotic results accurately reflect the finite-sample properties of LP and VAR procedures. The first set of simulations pertain to a simple ARMA(1,1) model, while the second set uses the popular structural macroeconomic model of Smets and Wouters (2007). Additional simulation results are reported in LABEL:app:sim_details.
5.1 Deriving the VARMA representation
We first show how we transform any given simulation DGP into the VARMA form assumed in Sections 3 and 4. The DGPs we consider have a vector MA representation
$y_{t}=C(L)u_{t},$ (5.1)
where $C(L)=I+\sum_{\ell=1}^{\infty}C_{\ell}L^{\ell}$, and $u_{t}$ is white noise with $\Sigma_{u}\equiv E(u_{t}u_{t}^{\prime})$. Given a choice of autoregressive lag length $p$, we seek to represent this process in VARMA($p,\infty$) form with MA component that is as small as possible. To this end, consider a linear least-squares projection of $y_{t}$ onto $p$ of its lags. We write the projection coefficients as $A_{\ell}(p)$, $\ell=1,2,\dots,p$, and define $A(L;p)\equiv I-\sum_{\ell=1}^{p}A_{\ell}(p)$. Letting $\tilde{C}(L;p)=\sum_{\ell=0}^{\infty}\tilde{C}_{\ell}(p)L^{\ell}\equiv A(L;p)C(L)$, equation (5.1) implies the VARMA($p,\infty$) representation
$y_{t}=\sum_{\ell=1}^{p}A_{\ell}(p)y_{t-\ell}+\sum_{\ell=0}^{\infty}\tilde{C}_{\ell}(p)u_{t-\ell}.$ (5.2)
Given a choice of $T$ as well as $\zeta$, it is straightforward to translate the representation given in (5.2) into our baseline form (3.1). [^9] Specifically, set $A_{\ell}=A_{\ell}(p)$, and recover $H$ and $D$ as the unique lower-triangular and diagonal matrices that satisfy $HDH^{\prime}=\tilde{C}_{0}(p)\Sigma_{u}\tilde{C}_{0}(p)^{\prime}$, with $1$’s along the main diagonal of $H$. Finally recover $\alpha_{\ell}=(HT^{-\zeta})^{-1}\tilde{C}_{\ell}(p)\tilde{C}_{0}(p)^{-1}H$ for $\ell=1,2,\dots$. We have thus arrived at a local-to-SVAR($p$) model that perfectly matches the second-moment properties of the original process $\{y_{t}\}$ in (5.1). The MA residual in (5.2) then measures the magnitude of the departure of the DGP from the best-fitting VAR($p$) model. As is the case for many structural dynamic macro models, the simulation DGPs we consider below cannot be represented exactly as finite-order VAR models. Our results will shed light on whether the resulting MA components are large enough to meaningfully distort VAR inference in practice.
5.2 Illustrative univariate model
We begin with an illustration based on a simple univariate ARMA(1,1) model. The purpose of this section is threefold: first, to get a sense of possible magnitudes for the degree of VAR misspecification $M$; second, to compare the actual with the worst-case lag polynomial $\alpha(L)$ in a simple and transparent environment; and third, to provide a visual illustration of our main theoretical results.
Model.
We consider an ARMA(1,1) process:
$y_{t}=\rho y_{t-1}+\varepsilon_{t}+\psi\varepsilon_{t-1},\quad\varepsilon_{t}\overset{i.i.d.}{\sim}N(0,\sigma^{2}).$ (5.3)
Throughout this section we set $\rho=0.9$, $\sigma^{2}=1$, and $\psi\in\{0,0.25,0.5,0.75\}$. The lagged MA term thus accounts for up to 36 per cent of the overall variance of the error term in (5.3).
@@ Results.
We first quantify the amount of misspecification. For a given lag length $p$ as well as reference values $T=240$ and $\zeta=1/2$, we can proceed as in Section 5.1 to translate the model (5.3) into an ARMA($p,\infty$) representation (5.2), with the AR parameters providing the best fit given $p$ [^10] Note that these best-fitting AR parameters differ from the parameter $\rho$ in (5.3) due to the serial correlation of the MA term. Table 5.1 shows the total amount of MA misspecification $M\equiv||\alpha(L)||$ as a function of $p$ and $\psi$. The table also reports the quantity $M^{2}/(1+M^{2})$, the minimax optimal ex ante model averaging weight on LP in 4.2. As expected, shorter lag lengths $p$ or larger MA coefficients $\psi$ mean greater misspecification, in the sense of larger $M$. We also see that, even for moderate $p$ and $\psi$, the degree of misspecification—and accordingly the optimal LP weight—can be large.
$M$ $\frac{M^{2}}{1+M^{2}}$
$p$ $\psi=0.25$ $\psi=0.50$ $\psi=0.75$ $\psi=0.25$ $\psi=0.50$ $\psi=0.75$
1 3.622 7.396 11.234 0.929 0.982 0.992
2 0.882 3.337 6.821 0.437 0.918 0.979
3 0.220 1.631 4.682 0.046 0.727 0.956
4 0.055 0.811 3.361 0.003 0.397 0.919
Table 5.1: $M$ and $\frac{M^{2}}{1+M^{2}}$ as a function of $p$ and $\psi$ in the ARMA(1,1) model (5.3), with the researcher estimating an AR($p$).
Next we show that, in this DGP, the implied misspecification polynomial $\alpha(L)$ can be very close to the theoretical least favorable polynomial $\alpha^{\dagger}(L;h)$ derived in Section 4.1, which represents the worst case for AR bias and coverage. Figure 5.1 shows $\alpha(L)$ (solid black) as well as $\alpha^{\dagger}(L;h)$ (colored, dashed and dotted) for horizons $h\in\{1,5,10\}$, and throughout setting $\psi=0.5$. For both $p=1$ as well as $p=4$, the actual MA polynomial $\alpha(L)$ implied by the ARMA(1,1) model (5.3) is very close—though not quite identical—to the worst-case $\alpha^{\dagger}(L;1)$ at horizon $h=1$. Hence, the least favorable lag polynomial is not some practically immaterial theoretical curiosity in the particular DGP considered here.
Figure 5.1: $\alpha(L)$ (black, solid) and least favorable $\alpha^{\dagger}(L)$ (colored, dashed and dotted) at various horizons for the simple ARMA(1,1) model (5.3), with the researcher estimating an AR($1$) (left) and an AR($4$) (right).
Finally, and consistent with our theory, we find that coverage can be poor for VAR confidence intervals, while LP intervals are robust to the presence of the MA term. Figure 5.2 reports coverage rates and median confidence interval lengths for the cases $\psi=0$ (no misspecification) and $\psi=0.25$ (moderate misspecification, with the lagged MA term accounting for around 6% of the variance of the error term). We throughout consider $p=1$, simulate 5,000 samples of size $T=240$, and then for each construct delta method as well as bootstrap LP and AR confidence intervals (assuming homoskedasticity), in blue and red, respectively. [^11] We construct percentile-$t$ bootstrap confidence intervals, with 5,000 bootstrap draws. The top panel reveals that, when the AR(1) model is in fact correctly specified (i.e., for $\psi=0$), then both LP and AR confidence intervals attain the nominal coverage probability of 90 per cent (left panel); furthermore, and also as expected, the AR confidence intervals are meaningfully shorter (right panel). In the misspecified case in the bottom panel, the AR confidence intervals instead substantially undercover, and particularly so at short horizons. This is consistent with our general theoretical results as well as our observation above that the actual misspecification lag polynomial $\alpha(L)$ is close to the horizon-$1$ worst-case one, $\alpha^{\dagger}(L;1)$. We also see that LP, on the other hand, exhibits at worst mild undercoverage.
AR(1) – Correct Specification
ARMA(1,1), $\psi=0.25$
Figure 5.2: Coverage probabilities (left) and median confidence interval length (right) for AR (red) and LP (blue) for delta method and bootstrap (solid, dashed, and dotted). The DGP is (5.3) with $\psi=0$ (top) and $\psi=0.25$ (bottom), with estimation lag length fixed at $p=1$.
5.3 Misspecification in Smets and Wouters (2007)
For our quantitative simulation exercise, we take as the DGP the well-known structural macroeconomic model of Smets and Wouters (2007). This environment is ideal for our purposes, as it allows us to closely mimic applied macroeconometric practice: the model is rich enough to match well the second-moment properties of standard macroeconomic time series data, yet at the same time features the interpretable, structural shocks typically studied in applied work, like monetary or cost-push shocks.
Framework.
We consider the model of Smets and Wouters, solved at its (quarterly) posterior mode parameterization. The econometrician observes the wage cost-push shock, inflation, wages, and total hours worked. The impulse response function of interest is that of inflation with respect to the cost-push shock. Our specification follows recent work that studies the role of labor markets in the post-2021 inflation spike (e.g., Bernanke and Blanchard, 2023). Results for alternative set-ups with other structural shocks and different shock identification schemes are reported in LABEL:app:sim_details.
Results.
We begin by quantifying the amount of misspecification, having represented the model in VARMA form (see Section 5.1) for $T=240$ and $\zeta=1/2$. Table 5.2 shows the total degree of misspecification $M$ as well as the minimax MSE-optimal weight $M^{2}/(1+M^{2})$ on LP in 4.2 as a function of the VAR lag length $p$. As anticipated, the larger $p$, the smaller $M$. Importantly, however, $M$ only declines extremely slowly with the lag length $p$. In particular, for lag lengths typical in applied practice (for quarterly data), misspecification is material, with $M\approx 3.23$ for a standard lag length of $p=4$, and $M\approx 1.89$ even for a long lag length of $p=8$. This suggests that—depending on the shape of $\alpha(L)$—there is potential for misspecification to materially affect VAR inference.
@@
Table 5.2: $M$ and $\frac{M^{2}}{1+M^{2}}$ as a function of $p$ in the structural model of Smets and Wouters (2007), with the researcher observing the cost-push shock, inflation, wages, and hours worked, and estimating a VAR($p$).
@@
Figure 5.3: See Figure 5.2. The DGP is the model of Smets and Wouters, and the researcher estimates the response of inflation to an observed cost-push shock. Lag length $p$ is selected using the AIC for the top panel, and set to $p=4$ and $p=8$ for the middle and bottom panels, respectively.
Figure 5.3 shows that VAR confidence intervals indeed severely undercover, while LP intervals remain robust. We simulate 5,000 samples of size $T=240$, and for each construct delta method as well as bootstrap LP and VAR confidence intervals (assuming homoskedasticity). The top panel displays our headline results, with lag length $p$ selected using the AIC. At all but very short horizons, VAR confidence intervals materially undercover, while LP throughout attains close to the nominal coverage level, consistent with our theoretical results. We emphasize that these results for VAR inference are obtained even though the lag length $p$ is selected using the AIC: the median selected lag length is $p=2$, which here is evidently insufficient to guard against material VAR bias and undercoverage. The middle and bottom panels then further illustrate our “no free lunch” results. For those panels, we instead manually select longer lag lengths: $p=4$ for the middle panel and $p=8$ for the bottom panel. VAR coverage is now closer to the nominal level for all horizons $h\leq p$ (consistent with 3.3), but at the same time confidence intervals become essentially as wide as for LP. At longer horizons we obtain the same picture as before: material undercoverage for VAR, and coverage close to the nominal level for LP. [^12] The VAR coverage distortion is not a small-sample phenomenon: LABEL:app:sim_details shows that similar coverage distortions arise with a larger sample size.
LABEL:app:sim_details shows that these conclusions extend to alternative shocks and alternative shock identification schemes. We there furthermore investigate what happens if the actual lag polynomial $\alpha(L)$ is replaced by the theoretical least favorable one. In that case, the magnitudes of VAR undercoverage are broadly comparable with those obtained under the actual $\alpha(L)$ implied by the Smets and Wouters (2007) model, revealing that the least favorable MA polynomial $\alpha^{\dagger}(L,\cdot)$ is not particularly pathological, echoing our earlier finding in the univariate ARMA(1,1) simulation DGP.
Summary.
The simulation results in this section suggest that our theoretical results are likely to have practical bite. In empirically realistic settings, LP confidence intervals have close-to-nominal coverage, while VAR misspecification is plausibly large enough (and similar enough to the worst-case misspecification) to cause severe VAR coverage distortions. A researcher may guard against such distortions by increasing the lag length beyond the values suggested by standard information criteria; however, in doing so, she will end up with confidence intervals of length comparable to those of LPs.

### v2

@@ 4.3 Bias-aware inference
We thus conclude that, while bias-aware VAR inference is possible in theory, in practice the gains relative to the simpler LP interval are small at best, unless we put an extremely tight bound on the noise-to-signal ratio.
4.4 Inference on multiple impulse responses
Since the least favorable MA polynomial derived in Section 4.2 depends on the horizon $h$ of interest, one might hope that VARs would not be as prone to bias and thereby undercoverage if interest centers on multiple impulse responses. Unfortunately, we now show that this is not the case by generalizing the worst-case bias formula to the multi-dimensional case and deriving the worst-case coverage of the Wald confidence ellipsoid.
Set-up.
We now consider inference on any combination of impulse responses for various horizons $h$, response variables $i^{*}$, and shocks $j^{*}$. When referring to impulse responses and estimators of these, we need to make the response variable and shock explicit in the notation. Thus, we write $\theta_{i^{*},j^{*},h,T}$, $\hat{\beta}_{i^{*},j^{*},h}$, and $\hat{\delta}_{i^{*},j^{*},h}$, with the definitions being the same as in Section 2. Let $k$ denote the total number of impulse responses of interest. We refer to the list of impulse responses by the collection of triples $\{(i_{a}^{*},j_{a}^{*},h_{a})\}_{a=1}^{k}$ indexing the response variable, shock variable, and horizon, respectively. Define the $k$-dimensional vectors of true impulse responses and LP and VAR estimators:
$\boldsymbol{\theta}_{T}\equiv\begin{pmatrix}\theta_{i_{1}^{*},j_{1}^{*},h_{1},T}\\ \vdots\\ \theta_{i_{k}^{*},j_{k}^{*},h_{k},T}\end{pmatrix},\quad\hat{\boldsymbol{\beta}}\equiv\begin{pmatrix}\hat{\beta}_{i_{1}^{*},j_{1}^{*},h_{1}}\\ \vdots\\ \hat{\beta}_{i_{k}^{*},j_{k}^{*},h_{k}}\end{pmatrix},\quad\hat{\boldsymbol{\delta}}\equiv\begin{pmatrix}\hat{\delta}_{i_{1}^{*},j_{1}^{*},h_{1}}\\ \vdots\\ \hat{\delta}_{i_{k}^{*},j_{k}^{*},h_{k}}\end{pmatrix}.$
It follows from Propositions 3.1 and 3.2 that
$\sqrt{T}\begin{pmatrix}\hat{\boldsymbol{\beta}}-\boldsymbol{\theta}_{T}\\ \hat{\boldsymbol{\delta}}-\boldsymbol{\theta}_{T}\end{pmatrix}\stackrel{{\scriptstyle d}}{{\to}}N\left(\begin{pmatrix}0_{k\times 1}\\ \operatorname{aBias}(\hat{\boldsymbol{\delta}})\end{pmatrix},\begin{pmatrix}\operatorname{aVar}(\hat{\boldsymbol{\beta}})&\operatorname{aCov}(\hat{\boldsymbol{\beta}},\hat{\boldsymbol{\delta}})\\ \operatorname{aCov}(\hat{\boldsymbol{\delta}},\hat{\boldsymbol{\beta}})&\operatorname{aVar}(\hat{\boldsymbol{\delta}})\end{pmatrix}\right),$ (4.2)
for a $k$-dimensional vector $\operatorname{aBias}(\hat{\boldsymbol{\delta}})$ (defined in the proof of Proposition 4.2 below) and $k\times k$ matrices $\operatorname{aVar}(\hat{\boldsymbol{\beta}})$, $\operatorname{aVar}(\hat{\boldsymbol{\delta}})$, and $\operatorname{aCov}(\hat{\boldsymbol{\beta}},\hat{\boldsymbol{\delta}})$ given in Corollary A.2 in Section A.3. This corollary also implies that the difference $\hat{\boldsymbol{\beta}}-\hat{\boldsymbol{\delta}}$ is asymptotically independent of $\hat{\boldsymbol{\delta}}$, which is not surprising given the general arguments of Hausman (1978) and the facts that (i) the asymptotic variances of the estimators are the same as in the model with $\alpha(L)=0$ and (ii) the VAR estimator is the quasi-MLE in such a model. It follows that $\operatorname{aVar}(\hat{\boldsymbol{\beta}})\geq\operatorname{aVar}(\hat{\boldsymbol{\delta}})$ in the positive semidefinite sense.
Worst-case bias.
The following result generalizes the univariate worst-case bias formula in Proposition 4.1.
Proposition 4.2.
Impose Assumption 2.1, with part (iii) holding for all shock indices $j_{1}^{*},\dots,j_{k}^{*}$, and let $\zeta=1/2$. Let $R$ be a constant matrix with $k$ rows. Then
$\max_{\alpha(L)\colon\|\alpha(L)\|\leq M}\|R\operatorname{aBias}(\hat{\boldsymbol{\delta}})\|^{2}=M^{2}\lambda_{\max}\!\left(R[\operatorname{aVar}(\hat{\boldsymbol{\beta}})-\operatorname{aVar}(\hat{\boldsymbol{\delta}})]R^{\prime}\right),$
where $\lambda_{\max}(B)$ denotes the largest eigenvalue of the matrix $B$.
Proof.
See Section B.3. ∎
The proposition shows that the worst-case squared norm of the bias of the VAR estimator $R\hat{\boldsymbol{\delta}}$ of $R\boldsymbol{\theta}_{T}$ is a function of two simple quantities: the bound $M$ on misspecification, and the largest eigenvalue of the difference $\operatorname{aVar}(R\hat{\boldsymbol{\beta}})-\operatorname{aVar}(R\hat{\boldsymbol{\delta}})$ between the variance-covariance matrices for the LP and VAR estimators. The latter eigenvalue equals $\max_{\|\varsigma\|=1}\{\operatorname{aVar}(\varsigma^{\prime}R\hat{\boldsymbol{\beta}})-\operatorname{aVar}(\varsigma^{\prime}R\hat{\boldsymbol{\delta}})\}$, i.e., the largest efficiency gain for VAR over LP across all linear combinations (with norm 1) of the estimated parameters. We thus see that there is still no free lunch: the worst-case bias is non-negligible if the VAR offers efficiency gains for any linear combination of the parameters of interest, echoing our univariate results. When $R$ is a row vector, then the proposition implies that our conclusions from Section 4.2 extend to inference on any linear combination of impulse responses (across horizons, variables, and/or shocks). In particular, the VAR confidence interval has fragile coverage even if the target parameter is the integral (i.e., sum) of impulse responses across multiple horizons, as in the fiscal multiplier applications reviewed in Ramey (2016).
Worst-case coverage of confidence ellipsoid.
We next derive the coverage of the conventional Wald confidence ellipsoid based on the VAR estimator. The level-($1-a$) confidence ellipsoid is given by
$\operatorname{CE}(\hat{\boldsymbol{\delta}})\equiv\left\{\tilde{\boldsymbol{\theta}}\in\mathbb{R}^{k}\colon T(\hat{\boldsymbol{\delta}}-\tilde{\boldsymbol{\theta}})^{\prime}\operatorname{aVar}(\hat{\boldsymbol{\delta}})^{-1}(\hat{\boldsymbol{\delta}}-\tilde{\boldsymbol{\theta}})\leq\chi_{1-a,k}^{2}\right\},$
where $\chi_{1-a,k}^{2}$ is the $1-a$ quantile of the $\chi^{2}$ distribution with $k$ degrees of freedom.
Corollary 4.6.
Impose Assumption 2.1, with part (iii) holding for all shock indices $j_{1}^{*},\dots,j_{k}^{*}$, and let $\zeta=1/2$. Assume also that $\operatorname{aVar}(\hat{\boldsymbol{\delta}})$ is non-singular. Then
$\min_{\alpha(L)\colon\|\alpha(L)\|\leq M}\lim_{T\to\infty}P(\boldsymbol{\theta}_{T}\in\operatorname{CE}(\hat{\boldsymbol{\delta}}))=F_{k}\left(\chi_{1-a,k}^{2};M^{2}\left[\lambda_{\max}(\operatorname{aVar}(\hat{\boldsymbol{\beta}})\operatorname{aVar}(\hat{\boldsymbol{\delta}})^{-1})-1\right]\right),$
where $F_{k}(x;c)$ is the cumulative distribution function, evaluated at point $x\geq 0$, of a non-central $\chi^{2}$ distribution with $k$ degrees of freedom and non-centrality parameter $c\geq 0$.
Proof.
See Section B.8. ∎
The worst-case coverage probability of the VAR confidence ellipsoid depends on three scalars: the bound $M$ on misspecification, the dimension $k$ of the ellipsoid, and the “multivariate relative standard error”
$\sqrt{\lambda_{\min}(\operatorname{aVar}(\hat{\boldsymbol{\delta}})\operatorname{aVar}(\hat{\boldsymbol{\beta}})^{-1})}=[\lambda_{\max}(\operatorname{aVar}(\hat{\boldsymbol{\beta}})\operatorname{aVar}(\hat{\boldsymbol{\delta}})^{-1})]^{-1/2}=\min_{\varsigma\in\mathbb{R}^{k}}\sqrt{\operatorname{aVar}(\varsigma^{\prime}\hat{\boldsymbol{\delta}})/\operatorname{aVar}(\varsigma^{\prime}\hat{\boldsymbol{\beta}})}.$
Again, the worst-case coverage distortion is an increasing function of the largest efficiency gain for VAR over LP across all linear combinations of the impulse responses. Since VAR impulse response estimates are often highly correlated across horizons, this suggests that the VAR undercoverage can in fact be particularly severe in the multivariate case.
Figures 4.4 and 4.5 show that the worst-case coverage probability of the confidence ellipsoid can be very poor even for small amounts of misspecification. The panels in the figures correspond to choices of the dimension $k\in\{2,5,10\}$ or the bound $M\in\{1,1.5,2\}$ on misspecification, respectively. Evidently, the coverage distortions can be severe for all of the $k$ considered in these plots even when $M=1$ (corresponding to a noise-to-signal ratio of $1/T$), especially if the multivariate relative standard error $\sqrt{\lambda_{\min}(\operatorname{aVar}(\hat{\boldsymbol{\delta}})\operatorname{aVar}(\hat{\boldsymbol{\beta}})^{-1})}\leq 1/4$. It further appears that the worst-case coverage distortion is monotonically decreasing in $k$, holding everything else constant; however, loosely speaking, the larger is $k$, the larger is the “chance” that there is some linear combination of the parameters for which VAR is much more efficient than LP, instead yielding a larger coverage distortion. The multivariate case thus overall echoes and reinforces our conclusions from the earlier univariate analysis.
Figure 4.4: Worst-case asymptotic coverage probability of the conventional 90% VAR confidence ellipsoid. Different panels: choices of the number $k$ of parameters in the ellipsoid. Horizontal axes: square root of smallest eigenvalue of the ratio of the VAR and LP variance-covariance matrices. Different lines: different bounds $M$ on $\|\alpha(L)\|$. The solid horizontal lines mark the nominal coverage probability $1-a=90\%$.
Figure 4.5: Worst-case asymptotic coverage probability of the conventional 90% VAR confidence ellipsoid. Different panels: different bounds $M$ on $\|\alpha(L)\|$. Horizontal axes: square root of smallest eigenvalue of the ratio of the VAR and LP variance-covariance matrices. Different lines: choices of the number $k$ of parameters in the ellipsoid. The solid horizontal lines mark the nominal coverage probability $1-a=90\%$.
@@ 5 Simulations
We now show that our asymptotic results accurately reflect the finite-sample properties of LP and VAR procedures. To do so, we consider a standard structural macroeconomic model as our DGP; in LABEL:app:simul_arma we present further illustrative results from a simple univariate model.
Framework.
Our DGP is the well-known structural macroeconomic model of Smets and Wouters (2007). This environment is ideal for our purposes, as it allows us to closely mimic applied macroeconometric practice: the model is rich enough to match quite well the second-moment properties of standard macroeconomic time series data, yet at the same time features the interpretable, structural shocks typically studied in applied work. We solve the model at the posterior mode estimated by Smets and Wouters. For our main exercise we will assume that the econometrician observes the wage cost-push shock, inflation, wages, and total hours worked. The impulse response function of interest is that of inflation with respect to the cost-push shock. In addition to being topical (e.g., see the recent work of Bernanke and Blanchard, 2023), this exercise is well-suited to cleanly illustrate the potential bite of our theoretical conclusions: wage cost-push shocks in the model of Smets and Wouters follow an ARMA($1$,$1$) process, and so short-lag VARs may be subject to material biases. Results for a monetary shock specification are reported in LABEL:app:sim_observedmp.
@@ Results.
Table 5.1: $M$ and $\frac{M^{2}}{1+M^{2}}$ as functions of $p$ in the model of Smets and Wouters, with the researcher observing the cost-push shock, inflation, wages, and hours worked, and estimating a VAR($p$).
@@
Figure 5.1: Coverage probabilities (left) and median confidence interval length (right) for VAR (red) and LP (blue) confidence intervals computed via the delta method or bootstrap (the latter are indicated with subscript “b” in the figure legends). DGP: Smets and Wouters (2007), cost-push shock. Lag length $p$ is selected using the AIC for the top panel, and fixed at $p=4$ and $p=8$ for the middle and bottom panels, respectively.
We begin by quantifying the amount of VAR misspecification in this DGP. Given a choice of lag length $p$, we represent the Smets and Wouters model in VARMA($p,\infty$) form (2.1) with $T=240$ and $\zeta=1/2$, where the VAR coefficients are selected to minimize the population sum of squared residuals.888 Let $y_{t}=C(L)u_{t}$ be the Wold decomposition of the observed data. Denote the linear least-squares projection coefficients of $y_{t}$ onto $p$ of its lags as $A_{\ell}(p)$, $\ell=1,2,\dots,p$, and define $A(L;p)\equiv I-\sum_{\ell=1}^{p}A_{\ell}(p)$. Letting $\tilde{C}(L;p)=\sum_{\ell=0}^{\infty}\tilde{C}_{\ell}(p)L^{\ell}\equiv A(L;p)C(L)$, we obtain a representation $y_{t}=\sum_{\ell=1}^{p}A_{\ell}(p)y_{t-\ell}+\sum_{\ell=0}^{\infty}\tilde{C}_{\ell}(p)u_{t-\ell}$. It is straightforward to orthogonalize the innovations in this representation and write it in the form (2.1), thus yielding the implied MA polynomial $\alpha(L)$. Table 5.1 shows, as a function of $p$, the total degree of misspecification $M=\|\alpha(L)\|$ as well as $M^{2}/(1+M^{2})$, the minimax optimal ex ante model averaging weight on LP in Corollary 4.2. As anticipated, the larger $p$, the smaller $M$. Importantly, however, $M$ only declines extremely slowly with the lag length $p$. In particular, for lag lengths typical in applied practice (for quarterly data), misspecification is material, with $M\approx 3.23$ for a standard lag length of $p=4$, and $M\approx 1.89$ even for a long lag length of $p=8$. This suggests that—depending on the shape of $\alpha(L)$—there is potential for misspecification to substantially affect VAR inference.
[…]
LABEL:app:sim_observedmp shows that these conclusions extend to monetary shock specifications. We furthermore show that the VAR coverage distortions in the actual Smets and Wouters DGP are comparable to those in a slightly perturbed DGP where we replace the model-implied MA lag polynomial $\alpha(L)$ by the theoretical least favorable one $\alpha^{\dagger}(L)$; thus, the latter is not particularly pathological, consistent with our discussion in Section 4.2.